# Quantum Hall Edges Beyond the Plasma Analogy

### v1

@@ III Mismatched edge correlations
Below we provide detailed examples of the difference between the angle coordinates (13) and (14). We begin with disk-shaped QH droplets whose confining potential is anisotropic, then consider anisotropic droplets obtained from the “edge-deformed potentials” introduced in [28], including square droplets where CG predictions are known analytically.
Anisotropic disks. Consider a QH droplet whose domain $\Omega$ is a disk, but whose confining potential is anisotropic as in Fig. 3(b) (so its level curves are not concentric circles). Edge modes then propagate with a velocity whose norm (2) is non-constant, leading to a canonical angle (3). When the droplet $\Omega$ is a disk, the arc length is $s=L\varphi/2\pi$ and the angle (3) becomes a non-trivial function $\theta(\varphi)$, expressed as a functional of the local norm $v(\varphi)$ of the edge velocity. The resulting map (13) is a non-trivial diffeomorphism of the circle since $\partial\Omega=S^{1}$.
@@
Now suppose one were to use the plasma analogy to predict edge correlations in this setup. Consider therefore a disk-shaped CG with uniform bulk density. The corresponding potential in the screening condition (8) is necessarily isotropic and the conformal map (11) is just the identity (up to rotation), so the coordinate $\alpha$ in the edge correlator (12) coincides with the polar angle, $\alpha=\varphi$. This manifestly differs from the canonical angle (3), exemplifying the announced mismatch.
@@
Then choose a radius $R$ for the QH droplet, and let the potential near the edge take the form
$V(r,\varphi)=V_{0}\Bigl{(}\pi R^{2}+\pi\bigl{(}r^{2}-R^{2}\bigr{)}\big{/}\theta^{\prime}(\varphi)\Bigr{)}$ (16)
for some monotonically increasing function $V_{0}(\cdot)$. The corresponding action-angle coordinates in a neighbourhood of the edge are
$\mathcal{A}=\pi R^{2}+\pi\bigl{(}r^{2}-R^{2}\bigr{)}\big{/}\theta^{\prime}(\varphi),\qquad\theta=\theta(\varphi).$ (17)
@@
A subclass of edge-deformed QH droplets is provided by flower deformations with an integer number $k$ of “petals”, for which the function $\theta(\varphi)$ in the potential (18) is given by
@@
Figure 4: Differences between coordinates $\theta(\varphi)$ and $\alpha(\varphi)$ for edge-deformed droplets (19) given by flower deformations (20) with $(k,\lambda)=(3,0.2)$ (purple), $(k,\lambda)=(3,0.4)$ (yellow), $(k,\lambda)=(2,0.4)$ (blue), and a square (21) (magenta). Equipotentials of the respective QH droplets are shown on the side.
Now suppose again one uses the plasma analogy to describe this situation. To this end, consider a CG with uniform bulk density, whose shape $\Omega$ is a flower droplet with a boundary given by Eq. (19) in polar coordinates, with some fixed area $\mathcal{A}$. The resulting edge density correlator takes the form (12) in terms of the restriction of the conformal map (11) to the boundary $\partial\Omega$. The problem is to find this map $\mapsto G(\textbf{x})$ x ). As before, $G(\textbf{x})$ is uniquely fixed by the shape of the droplet through the Riemann mapping theorem. It admits no known analytical expression (unless $k=2$), but it can be built numerically using Fekete points [49]. This is how Fig. 4 was obtained: in it, we show the difference $\theta(\varphi)-\alpha(\varphi)$ for flower droplets with $k=2$ or $3$ petals. The difference is manifestly non-constant when $k=3$, confirming that $\theta(\varphi)$ and $\alpha(\varphi)$ differ by more than a rotation. In fact, the mismatch worsens when $\lambda$ increases. Similar conclusions hold for all droplets with $k\neq 2$ petals, and generally for the vast majority of QH droplets confined by edge-deformed potentials (18).
The exception $k=2$ corresponds to droplets whose boundary (19) is an ellipse. In that case, the conformal map (11) can be found analytically [34] and the CG edge correlator (12) coincides with the correlator (6) of a QH droplet in the potential (18); see the straight line where $\theta(\varphi)=\alpha(\varphi)$ in Fig. 4. There is a deep reason for this coincidence: the distribution of particles in an elliptic CG coincides with the distribution of eigenvalues of random matrices in the elliptic Ginibre ensemble [17, 9], which in turn coincides with the density of LLL wave functions in a harmonic potential.
A geometric way to understand why $k=2$ is exceptional is to bluntly impose $\theta(\varphi)=\alpha(\varphi)$ on the edge (19), then attempt to view $\alpha(\varphi)$ as the boundary value of a bijective conformal map (11). One readily verifies that this fails for any flower deformation (20), unless $k=2$. The failure is due to zeros or poles of the would-be conformal extension, which always appear outside $\Omega$ unless $k=2$; see the SM [34] for details. More generally, for any edge-deformed droplet, imposing $\theta(\varphi)=\alpha(\varphi)$ on the curve (19) makes it impossible to view $\alpha(\varphi)$ as the restriction of a bijective conformal map (11), except in exceedingly fine-tuned cases. This underscores again that the plasma analogy is typically unreliable on the edge of anisotropic QH droplets.
@@
in terms of the polar angle $\varphi$ and the map $G$ in (11), which is holomorphic in the complex coordinate $x+\mathrm{i}y$. One can again compare the QH and CG angles (21)–(22): their difference $\theta(\varphi)-\alpha(\varphi)$, shown in Fig. 4, fails to be a constant, exhibiting once more the QH/CG mismatch.
@@ IV Consequences for
Differences between QH droplets and their would-be plasma analogues affect the computation of physical quantities. Indeed, the mismatch between the correlators (6) and (12) entails different predictions for the contribution of low-energy edge modes to, say, linear-response coefficients of anisotropic droplets. Here we demonstrate its consequences for fluctuations of local operators, known as linear statistics in the CG literature [51, 12, 52]. As we shall explain, the fluctuation-dissipation theorem makes it possible to probe such fluctuations through microwave absorption experiments that are within reach using current technologies [29, 30, 31, 32, 33].
Fluctuations in QH droplets. To begin, consider the QH side. Given a smooth real function $U(\textbf{x})$, define the Fock-space operator
@@
where $\hat{c}^{(\dagger)}(\textbf{x})$ are local fermionic annihilation (creation) operators, $\{\hat{c}(\textbf{x}_{1}),\hat{c}^{\dagger}(\textbf{x}_{2})\}=\delta^{(2)}(\textbf{x}_{1}-\textbf{x}_{2})$. We wish to evaluate the variance of $\hat{U}$ in the $\nu=1$ QH ground state. By Wick’s theorem, one can express the variance as
@@
where $C(\textbf{x},\textbf{y})$ is the electronic two-point correlation function. The latter is short ranged in the bulk, where
$|C(\textbf{x},\textbf{y})|_{\textrm{bulk}}\sim\frac{1}{2\pi\ell^{2}}\mathrm{e}^{-\frac{|\textbf{x}-\textbf{y}|^{2}}{4\ell^{2}}}$ (25)
with $\ell$ the magnetic length introduced below (5). By contrast, correlations are long ranged along the edge, as mentioned around Eq. (6). A more precise statement, derived in [28] for edge-deformed potentials, is
@@
Fluctuations in CGs. On the CG side, the plasma analogue of the quantum variance (24) is the thermal fluctuation of the linear statistics $\sum_{j=1}^{N}U(\textbf{x}_{j})$, where $U$ is the same function as in Eq. (23). One is usually interested in the thermodynamic limit,
$\operatorname{Var}[U]_{\textrm{CG}}\equiv\lim_{N\to\infty}\operatorname{Var}\biggl{[}\sum_{j=1}^{N}U(\textbf{x}_{j})\biggr{]}.$ (28)
@@
where $U_{\alpha}$ is defined analogously to $U_{\theta}$ in (27), but using the conformal coordinate (14) instead of the canonical one (13). For the free-fermion value $\beta=2$, Eq. (29) looks identical to Eq. (27), up to one key difference: the angles $\theta,\alpha$ stem from different maps—(5) versus (11)—, leading to distinct functions $U_{\theta}$, $U_{\alpha}$.
Mismatch. Returning to the examples of Sec. III, it is straightforward to illustrate the quantitative mismatch between the fluctuations (27) and (29). Indeed, take for simplicity the function $U(x,y)=x$ for which $|\nabla U|^{2}=1$, so that the bulk terms in Eqs. (27) and (29) are proportional to the droplet’s area. The resulting total variances are plotted in Fig. 5 for flowers with $k=3$ petals in Eqs. (19)–(20), showing that QH and CG predictions disagree and that the mismatch worsens as the anisotropy increases. The same conclusions can be drawn for the square droplet: in that case, the variances (27) and (29) turn out to be
@@
which manifestly differ. Their exact values can be expressed using an infinite series and a hypergeometric function, respectively [34].
Figure 5: The total variance of $U(x,y)=x$ for flower droplets with $k=3$ petals and $\lambda$ ranging from 00 to $0.4$. Main figure: The variance (27) for QH droplets is compared to the CG result (29) obtained from a numerical evaluation of the conformal map (11), including an extrapolation based on Monte-Carlo (MC) data for the CG at $\lambda=0,0.1,0.2,0.3,0.4$ [34]. The two match for $\lambda=0$ but not $\lambda>0$. Inset: Details of the extrapolation for $\lambda=0.3$. The variance is extrapolated by a fit of the form $a+b/\sqrt{N}$ for $N=64,128,256$ particles. The intercept with the $y$ axis yields the coefficient $a$, which is shown as an empty green circle and differs significantly from its QH counterpart. To further illustrate this, we include a Monte-Carlo point for $N=1024$, which has a larger error bar but remains consistent with a leading $1/\sqrt{N}$ correction.
@@
Note that the link (32) between absorption and variance was to be expected, given the fluctuation-dissipation theorem. In fact, the latter ensures that the full variance (27)—including the bulk term—is the integral of the full absorption rate—including effects of higher Landau levels. No such bulk term appears in Eq. (32) since the rate (31) is limited to low-energy transitions within the LLL.
V Conclusion
In this paper, the plasma analogy was tested on QH droplets formed by anisotropic confining potentials. While it reproduces QH data in the bulk, the analogy generally fails at the edge—apart from exceptional, fine-tuned scenarios. Such discrepancies are expected to be directly relevant for experiments, both because of advances in measurement techniques for solid-state systems [55, 56, 32, 33] and increases in the precision of quantum simulators [21, 22, 23, 24, 25, 26, 27]. In each case, deviations from leading-order results are becoming visible, which is where the conclusions of this work are expected to be most striking.
We showed that the edge discrepancies stem from an elementary mismatch between QH and CG physics: the former is incompressible hydrodynamics, the latter is electrostatics. In geometric terms, QH is symplectic while CG is conformal. This distinction entails different angular dependencies of the corresponding correlators (6) and (12), which in turn affects e.g. fluctuations of local observables—quantities proposed to be measurable in the near future [29, 30, 31, 32, 33]. Such a breakdown of the plasma analogy is similarly expected near the edge of fractional QH states, as electron interactions are unlikely to improve the situation.
@@ Acknowledgments
We are grateful to Alexander Abanov, Yacin Ameur, Mathieu Beauvillain, Gabriel Cardoso, Laurent Charles, Benoit Douçot, Nathan Goldman, Gian Michele Graf, Hans Hansson, Leonardo Mazza, Gaultier Lambert, Alberto Nardin, Marios Petropoulos, Nicolas Rougerie, Gregory Schehr, Ood Shabtai and Paul Wiegmann for fruitful discussions on closely related subjects. P.M. gratefully acknowledges financial support from the Wenner-Gren Foundations (Grant No. FT2022-0002). B.L. acknowledges financial support from the Swiss National Science Foundation (Postdoc.Mobility Grant No. 214461).
@@
This supplemental material has three parts. Parts A and B respectively review the physics of quantum Hall (QH) droplets and Coulomb gases (CGs). Part C complements the main text by showing that the parametrization of the edge by a canonical angle coordinate generally fails to extend to the biholomorphism required in the CG setting, despite the fact that local conformal maps do reproduce the correlations of edge-deformed QH droplets.
@@
Here we review certain aspects of (anisotropic) QH droplets used in the main text. The presentation is limited to droplets of free fermions, although a similar picture is expected for fractional QH states as well. The key formulas are Eqs. (40)–(42), which are repeatedly used in the main text. We refer to [28] for more details.
@@
labeled by an angular momentum quantum number $m=0,1,2,\ldots$, where $z\equiv(x+\mathrm{i}y)/\sqrt{2}\ell$ is a dimensionless complex coordinate. The wave function of any state in the LLL is thus a holomorphic function times the Gaussian factor $\mathrm{e}^{-|z|^{2}/2}$.
We denote by $P=\sum_{m=0}^{\infty}|\phi_{m}\rangle\langle\phi_{m}|$ the LLL projection operator. One of its key properties is to make space non-commutative: $[PxP,PyP]=\mathrm{i}\ell^{2}$. Thus, in the LLL, space becomes akin to a symplectic manifold and the magnetic length (34) is analogous to Planck’s constant. We therefore refer to the limit of strong magnetic fields ($\ell\to 0$) as a “classical limit.”
@@
This is written here in terms of the canonical angle $\theta$ on the equipotential $\mathcal{A}_{m}=2\pi m\ell^{2}$, the distance $d$ from x to the equipotential, the constant angular velocity $\dot{\theta}(\mathcal{A}_{m})=\tfrac{2\pi}{qB}V_{0}^{\prime}(\mathcal{A}_{m})$ as in Eq. (4), and the Euclidean norm (2) of the guiding-center drift velocity on the equipotential. The classical phase $m\theta$ gives all the winding of $\psi_{m}$ with $\theta$, and is accompanied by a slowly varying quantum correction $\Theta_{m}(\textbf{x})$ that we leave unspecified here. Note the intuition in Eq. (37): it confirms that the probability density $|\psi_{m}(\textbf{x})|^{2}$ is localized as a Gaussian with width $\ell$ on the $m^{\text{th}}$ equipotential of $V$, and that the corresponding probability density $|\psi_{m}(\textbf{x})|^{2}$ is proportional to the inverse of the guiding-center velocity. This agrees with expectations based on classical reasoning: the link between quantum wave function and classical local velocity is one of the hallmarks of the WKB method.
@@
which is similarly found to be short ranged in the bulk [recall Eq. (25)] but long ranged along the edge. Specifically, plugging the wave functions (37) into Eq. (39) and summing over $k\equiv N-m$ yields a geometric series due to the leading phase $\mathrm{e}^{\mathrm{i}m\theta}$. This results in the following asymptotic expression, verified in [28] for edge-deformed droplets:
@@
Here $\Theta(\textbf{x}_{1},\textbf{x}_{2})$ is a complicated overall phase, $\omega_{\textrm{F}}\equiv\dot{\theta}(\mathcal{A}_{N})$ is the angular Fermi velocity, $v(\theta)\equiv v_{N}(\theta)$ is the norm (2) of the edge velocity, and we assume $\theta_{\textbf{x}}\neq\theta_{\textbf{y}}$. We have thus recovered Eq. (26) in the main text, with the normalization $\mathcal{N}\equiv\omega_{\textrm{F}}/2\pi^{3/2}\ell$. Finally, the connected density-density correlator for free fermions is
@@
$\rho(\textbf{x})\equiv\frac{1}{N}\Big{<}\sum_{j=1}^{N}\delta(\textbf{x}-\textbf{x}_{j})\Big{>}.$ (46)
This is normalized so that $\int\mathrm{d}\textbf{x}\,\rho(\textbf{x})=1$, since the total number $N$ of particles does not fluctuate. For an isotropic harmonic potential $W(\textbf{x})=|\textbf{x}|^{2}$, the density (46) in the thermodynamic limit is uniform in a disk and zero outside. In what follows, we explain this fact and various others, e.g. involving the density-density correlator $\big{<}\sum_{i,j}\delta(\textbf{x}-\textbf{x}_{i})\delta(\textbf{x}-\textbf{x}_{j})\big{>}$ or the variance of the one-body observable $\sum_{j=1}^{N}U(\textbf{x}_{j})$ for some (smooth) function $U$.
@@
Let us ask how the mean density (46) behaves in the limit $N\to\infty$. In statistical mechanics, there is usually a non-trivial competition between energy and entropy, but this is not so here (at least to leading order) due to long-range Coulomb interactions: the energy (43) is of order $N^{2}$, while entropic fluctuations turn out to be of order $N$ [43]. The density profile can thus be found from mere energy minimization, without thermal effects. For large $N$, one can approximate sums by integrals to rewrite the energy (43) as $E\approx(N^{2}/2)\mathcal{E}[\mu]$, where what one seeks to minimize is the energy functional
@@
The minimizer $\mu=\rho$ is to be found among all possible densities $\mu$, i.e. non-negative functions that integrate to one, $\int\mathrm{d}\textbf{x}\,\mu(\textbf{x})=1$. Under mild technical conditions on the potential $W$, one can show that the minimizer exists, is unique, and has compact support; see [49] for a mathematical discussion. The region where $\rho$ is supported is identified as the droplet $\Omega$; then $\rho(\textbf{x})=0$ for all $\textbf{x}\notin\Omega$.

### v2

@@ III Mismatched edge correlations
Below, we provide detailed examples of the difference between the angle coordinates in Eqs. (13) and (14). We begin with disk-shaped QH droplets whose confining potential is anisotropic, then consider anisotropic droplets of the form introduced in [28], including square droplets where CG predictions are known analytically.
Anisotropic disks. Consider a QH droplet whose domain $\Omega$ is a disk but whose confining potential is anisotropic as in Fig. 3(b) (so its level curves are not concentric circles). Edge modes then propagate with a velocity whose norm (2) is nonconstant, leading to a canonical angle (3). When the droplet $\Omega$ is a disk, the arc length is $s=L\varphi/2\pi$, and the angle (3) becomes a nontrivial function $\theta(\varphi)$, expressed as a functional of the local norm $v(\varphi)$ of the edge velocity. The resulting map (13) is a nontrivial diffeomorphism of the circle since $\partial\Omega=S^{1}$.
@@
Now suppose one were to use the plasma analogy to predict edge correlations in this setup. Consider therefore a disk-shaped CG with uniform bulk density. The corresponding potential in the screening condition (8) is necessarily isotropic, and the conformal map (11) is just the identity (up to rotation), so the coordinate $\alpha$ in the edge correlator (12) coincides with the polar angle: $\alpha=\varphi$. This manifestly differs from the canonical angle (3), exemplifying the announced mismatch.
@@
Then, choose a radius $R$ for the QH droplet, and let the potential near the edge take the form
$V(r,\varphi)=V_{0}\bigl{(}\pi R^{2}+\pi\bigl{[}r^{2}-R^{2}\bigr{]}\big{/}\theta^{\prime}(\varphi)\bigr{)}$ (16)
for some monotonically increasing function $V_{0}(\cdot)$. The corresponding action-angle coordinates in a neighborhood of the edge are
$\mathcal{A}=\pi R^{2}+\pi\bigl{[}r^{2}-R^{2}\bigr{]}\big{/}\theta^{\prime}(\varphi),\qquad\theta=\theta(\varphi).$ (17)
@@
A subclass of edge-deformed QH droplets is provided by flower deformations with an integer number $k$ of petals, for which the function $\theta(\varphi)$ in the potential (18) is given by
@@
Figure 4: Differences between coordinates $\theta(\varphi)$ and $\alpha(\varphi)$ for edge-deformed droplets (19) given by flower deformations (20), with $(k,\lambda)=(3,0.2)$ (purple), $(k,\lambda)=(3,0.4)$ (yellow), $(k,\lambda)=(2,0.4)$ (blue), and a square (21) (magenta). Equipotentials of the respective QH droplets are shown on the side.
Now suppose again one uses the plasma analogy to describe this situation. To this end, consider a CG with uniform bulk density, whose shape $\Omega$ is a flower droplet with a boundary given by Eq. (19) in polar coordinates, with some fixed area $\mathcal{A}$. The resulting edge density correlator takes the form (12) in terms of the restriction of the conformal map (11) to the boundary $\partial\Omega$. The problem is to find this map $\mapsto G(\textbf{x})$ x ). As before, $G(\textbf{x})$ is uniquely fixed by the shape of the droplet through the Riemann mapping theorem. It admits no known analytical expression (unless $k=2$), but it can be built numerically using Fekete points [52]. This is how Fig. 4 was obtained: in it, we show the difference $\theta(\varphi)-\alpha(\varphi)$ for flower droplets with $k=2$ or $3$ petals. The difference is manifestly nonconstant when $k=3$, confirming that $\theta(\varphi)$ and $\alpha(\varphi)$ differ by more than a rotation. In fact, the mismatch worsens when $\lambda$ increases. Similar conclusions hold for all droplets with $k\neq 2$ petals and generally for the vast majority of QH droplets confined by edge-deformed potentials (18).
The exception $k=2$ corresponds to droplets whose boundary, given by Eq. (19), is an ellipse. In that case, the conformal map (11) can be found analytically [34], and the CG edge correlator (12) coincides with the correlator (6) of a QH droplet in the potential (18); see the straight line where $\theta(\varphi)=\alpha(\varphi)$ in Fig. 4. There is a reason for this coincidence: the distribution of particles in an elliptic CG coincides with the distribution of eigenvalues of random matrices in the elliptic Ginibre ensemble [17, 9], which in turn coincides with the density of LLL wave functions in a harmonic potential.
A geometric way to understand why $k=2$ is exceptional is to bluntly impose $\theta(\varphi)=\alpha(\varphi)$ on the edge given by Eq. (19), then attempt to view $\alpha(\varphi)$ as the boundary value of a bijective conformal map (11). One readily verifies that this fails for any flower deformation (20), unless $k=2$. The failure is due to zeros or poles of the would-be conformal extension, which always appear outside $\Omega$ unless $k=2$; see the SM [34] for details. More generally, for any edge-deformed droplet, imposing $\theta(\varphi)=\alpha(\varphi)$ on the curve (19) makes it impossible to view $\alpha(\varphi)$ as the restriction of a bijective conformal map (11), except in exceedingly fine-tuned cases. This underscores again that the plasma analogy is typically unreliable on the edge of anisotropic QH droplets.
@@
in terms of the polar angle $\varphi$ and the map $G$ in Eq. (11), which is holomorphic in the complex coordinate $x+\mathrm{i}y$. One can again compare the QH and CG angles in Eqs. (21) and (22): their difference $\theta(\varphi)-\alpha(\varphi)$, shown in Fig. 4, fails to be a constant, exhibiting once more the QH/CG mismatch.
@@ IV Consequences for
Differences between QH droplets and their would-be plasma analogues affect the computation of physical quantities. Indeed, the mismatch between the correlators in Eqs. (6) and (12) entails different predictions for the contribution of low-energy edge modes to, say, linear-response coefficients of anisotropic droplets. Here, we demonstrate its consequences for fluctuations of local operators, known as linear statistics in the CG literature [54, 12, 55]. As we shall explain, the fluctuation-dissipation theorem makes it possible to probe such fluctuations through microwave absorption experiments that are within reach using current technologies [29, 30, 31, 32, 33].
Fluctuations in quantum Hall droplets. To begin, consider the QH side. Given a smooth real function $U(\textbf{x})$, define the Fock-space operator
@@
where $\hat{c}^{(\dagger)}(\textbf{x})$ are local fermionic annihilation (creation) operators satisfying $\{\hat{c}(\textbf{x}_{1}),\hat{c}^{\dagger}(\textbf{x}_{2})\}=\delta^{(2)}(\textbf{x}_{1}-\textbf{x}_{2})$. We wish to evaluate the variance of $\hat{U}$ in the $\nu=1$ QH ground state. By Wick’s theorem, one can express the variance as
@@
where $C(\textbf{x},\textbf{y})$ is the electronic two-point correlation function. The latter is short-ranged in the bulk, where
$|C(\textbf{x},\textbf{y})|_{\textrm{bulk}}\sim\frac{1}{2\pi\ell^{2}}\exp\biggl{(}-\frac{|\textbf{x}-\textbf{y}|^{2}}{4\ell^{2}}\biggr{)}$ (25)
with $\ell$ the magnetic length introduced below Eq. (5). By contrast, correlations are long-ranged along the edge, as mentioned around Eq. (6). A more precise statement, derived in [28] for edge-deformed potentials, is
@@
Fluctuations in Coulomb gases. On the CG side, the plasma analogue of the quantum variance (24) is the thermal fluctuation of the linear statistics $\sum_{j=1}^{N}U(\textbf{x}_{j})$, where $U$ is the same function as in Eq. (23). One is usually interested in the thermodynamic limit,
$\operatorname{Var}[U]_{\textrm{CG}}\equiv\lim_{N\to\infty}\operatorname{Var}\!\left[\sum_{j=1}^{N}U(\textbf{x}_{j})\right]\!.$ (28)
@@
where $U_{\alpha}$ is defined analogously to $U_{\theta}$ in Eq. (27), but using the conformal coordinate (14) instead of the canonical one (13). For the free-fermion value $\beta=2$, Eq. (29) looks identical to Eq. (27), up to one key difference: the angles $\theta,\alpha$ stem from different maps—(5) vs (11)—and thus lead to distinct functions $U_{\theta}$, $U_{\alpha}$.
Mismatch. Returning to the examples of Sec. III, it is straightforward to illustrate the quantitative mismatch between the fluctuations in Eqs. (27) and (29). Indeed, take for simplicity the function $U(x,y)=x$ for which $|\nabla U|^{2}=1$, so that the bulk terms in Eqs. (27) and (29) are proportional to the droplet’s area. The resulting total variances are plotted in Fig. 5 for flowers with $k=3$ petals in Eqs. (19)–(20), showing that QH and CG predictions disagree and that the mismatch worsens as the anisotropy increases. The same conclusions can be drawn for the square droplet: in that case, the variances in Eqs. (27) and (29) turn out to be
@@
which manifestly differ. Their exact values can be expressed using an infinite series and a generalized hypergeometric function, respectively [34].
Figure 5: The total variance of $U(x,y)=x$ for flower droplets with $k=3$ petals and $\lambda$ ranging from 00 to $0.4$. Main figure: The variance (27) for QH droplets is compared with the CG result (29) obtained from a numerical evaluation of the conformal map (11), including an extrapolation based on Monte Carlo (MC) data for the CG at $\lambda=0$, $0.1$, $0.2$, $0.3$, and $0.4$ [34]. The two match for $\lambda=0$ but not $\lambda>0$. Inset: Details of the extrapolation for $\lambda=0.3$. The variance is extrapolated by a fit of the form $a+b/\sqrt{N}$ for $N=64$, $128$, and $256$ particles. The intercept with the $y$ axis yields the coefficient $a$, which is shown as an empty green circle and differs significantly from its QH counterpart. To further illustrate this, we include an MC point for $N=1024$, which has a larger error bar but remains consistent with a leading $1/\sqrt{N}$ correction.
@@
Note that the link in Eq. (32) between absorption and variance was to be expected, given the fluctuation-dissipation theorem. In fact, the latter ensures that the full variance (27)—including the bulk term—is the integral of the full absorption rate, including effects of higher Landau levels. No such bulk term appears in Eq. (32) since the rate (31) is limited to low-energy transitions within the LLL.
V Conclusions
In this paper, the plasma analogy was tested on integer QH droplets formed by anisotropic confining potentials. While it reproduces QH data in the bulk, the analogy generally fails at the edge—apart from exceptional, fine-tuned scenarios. Such discrepancies are expected to be directly relevant for experiments, both because of advances in measurement techniques for solid-state systems [59, 60, 32, 33] and increases in the precision of quantum simulators [21, 22, 23, 24, 25, 26, 27]. In each case, deviations from leading-order results are becoming visible, which is where the conclusions of this paper are expected to be most striking.
We showed that the edge discrepancies stem from an elementary mismatch between QH and CG physics: the former is incompressible hydrodynamics, the latter is electrostatics. In geometric terms, QH is symplectic, while CG is conformal. This distinction entails different angular dependencies of the corresponding correlators in Eqs. (6) and (12), which in turn affects e.g. fluctuations of local observables—quantities proposed to be measurable in the near future [29, 30, 31, 32, 33]. Such a breakdown of the plasma analogy is similarly expected near the edge of fractional QH states, as electron interactions are unlikely to improve the situation. It would be interesting to study this question for fractional droplets and other related systems where interactions play an important role, such as quantum dots or artificial atoms [61, 62, 63] and QH droplets with quasihole excitations [64].
@@ Acknowledgments
We are grateful to Alexander Abanov, Yacin Ameur, Mathieu Beauvillain, Gabriel Cardoso, Laurent Charles, Benoit Douçot, Nathan Goldman, Gian Michele Graf, Hans Hansson, Leonardo Mazza, Gaultier Lambert, Alberto Nardin, Marios Petropoulos, Nicolas Rougerie, Gregory Schehr, Ood Shabtai, and Paul Wiegmann for fruitful discussions on closely related subjects. P.M. acknowledges financial support from the Wenner-Gren Foundations (Grant No. FT2022-0002). B.L. acknowledges financial support from the Swiss National Science Foundation (Postdoc.Mobility Grant No. 214461).
@@
This supplemental material has three parts. Parts A and B respectively review the physics of quantum Hall (QH) droplets and Coulomb gases (CGs). Part C complements the main text by showing that the parameterization of the edge by a canonical angle coordinate generally fails to extend to the biholomorphism required in the CG setting, despite the fact that local conformal maps do reproduce the correlations of edge-deformed QH droplets.
@@
Here, we review certain aspects of (anisotropic) QH droplets used in the main text. The presentation is limited to droplets of free fermions, although a similar picture is expected for fractional QH states as well. The key formulas are Eqs. (40)–(42), which are repeatedly used in the main text. We refer to [28] for more details.
@@
labeled by an angular momentum quantum number $m=0,1,2,\ldots$, where $z\equiv(x+\mathrm{i}y)/\sqrt{2}\ell$ is a dimensionless complex coordinate. The wave function of any state in the LLL is thus a holomorphic function times the Gaussian factor $\mathrm{e}^{-|z|^{2}/2}$. For weak harmonic potentials, the story dates back to [37, 38] and is closely related to the case $V=0$; see e.g. [28].
We denote by $P=\sum_{m=0}^{\infty}|\phi_{m}\rangle\langle\phi_{m}|$ the LLL projection operator. One of its key properties is to make space noncommutative: $[PxP,PyP]=\mathrm{i}\ell^{2}$. Thus, in the LLL, space becomes akin to a symplectic manifold and the magnetic length (34) is analogous to Planck’s constant. We therefore refer to the limit of strong magnetic fields ($\ell\to 0$) as a classical limit.
@@
This is written here in terms of the canonical angle $\theta$ on the equipotential $\mathcal{A}_{m}=2\pi m\ell^{2}$, the distance $d$ from x to the equipotential, the constant angular velocity $\dot{\theta}(\mathcal{A}_{m})=\tfrac{2\pi}{qB}V_{0}^{\prime}(\mathcal{A}_{m})$ as in Eq. (4), and the Euclidean norm (2) of the guiding-center drift velocity on the equipotential. The classical phase $m\theta$ gives all the winding of $\psi_{m}$ with $\theta$, and is accompanied by a slowly varying quantum correction $\Theta_{m}(\textbf{x})$ that we leave unspecified here. Note the intuition in Eq. (37): it confirms that the probability density $|\psi_{m}(\textbf{x})|^{2}$ is localized as a Gaussian with width $\ell$ on the $m^{\text{th}}$ equipotential of $V$ and that the corresponding probability density $|\psi_{m}(\textbf{x})|^{2}$ is proportional to the inverse of the guiding-center velocity. This agrees with expectations based on classical reasoning: the link between quantum wave function and classical local velocity is one of the hallmarks of the WKB method.
@@
which is similarly found to be short-ranged in the bulk [recall Eq. (25)] but long-ranged along the edge. Specifically, plugging the wave functions (37) into Eq. (39) and summing over $k\equiv N-m$ yields a geometric series due to the leading phase $\mathrm{e}^{\mathrm{i}m\theta}$. This results in the following asymptotic expression, verified in [28] for edge-deformed droplets:
@@
Here, $\Theta(\textbf{x}_{1},\textbf{x}_{2})$ is a complicated overall phase, $\omega_{\textrm{F}}\equiv\dot{\theta}(\mathcal{A}_{N})$ is the angular Fermi velocity, $v(\theta)\equiv v_{N}(\theta)$ is the norm (2) of the edge velocity, and we assume $\theta_{\textbf{x}}\neq\theta_{\textbf{y}}$. We have thus recovered Eq. (26) in the main text, with the normalization $\mathcal{N}\equiv\omega_{\textrm{F}}/2\pi^{3/2}\ell$. Finally, the connected density-density correlator for free fermions is
@@
$\rho(\textbf{x})\equiv\frac{1}{N}\!\left\langle\sum_{j=1}^{N}\delta(\textbf{x}-\textbf{x}_{j})\right\rangle\!.$ (46)
This is normalized so that $\int\mathrm{d}\textbf{x}\,\rho(\textbf{x})=1$, since the total number $N$ of particles does not fluctuate. For an isotropic harmonic potential $W(\textbf{x})=|\textbf{x}|^{2}$, the density (46) in the thermodynamic limit is uniform in a disk and zero outside. In what follows, we explain this fact and various others, e.g. involving the density-density correlator $\bigl{\langle}\sum_{i,j}\delta(\textbf{x}-\textbf{x}_{i})\delta(\textbf{x}-\textbf{x}_{j})\bigr{\rangle}$ or the variance of the one-body observable $\sum_{j=1}^{N}U(\textbf{x}_{j})$ for some (smooth) function $U$.
@@
Let us ask how the mean density (46) behaves in the limit $N\to\infty$. In statistical mechanics, there is usually a nontrivial competition between energy and entropy, but this is not so here (at least to leading order) due to long-range Coulomb interactions: the energy (43) is of order $N^{2}$, while entropic fluctuations turn out to be of order $N$ [45]. The density profile can thus be found from mere energy minimization, without thermal effects. For large $N$, one can approximate sums by integrals to rewrite the energy (43) as $E\approx(N^{2}/2)\mathcal{E}[\mu]$, where what one seeks to minimize is the energy functional
@@
The minimizer $\mu=\rho$ is to be found among all possible densities $\mu$, i.e. nonnegative functions that integrate to one, $\int\mathrm{d}\textbf{x}\,\mu(\textbf{x})=1$. Under mild technical conditions on the potential $W$, one can show that the minimizer exists, is unique, and has compact support; see [52] for a mathematical discussion. The region where $\rho$ is supported is identified as the droplet $\Omega$; then $\rho(\textbf{x})=0$ for all $\textbf{x}\notin\Omega$.